# Continuous Vision-Language-Action Co-Learning with Semantic-Physical Alignment for Behavioral Cloning

### v3

@@ Introduction
  • We propose a multimodal continuous co-learning mechanism integrating vision, language, and proprioceptive to model temporal dependencies in latent space, ensuring smooth and consistent transitions across action states.
@@
LCM leverages BC methods to infer executable behaviors from heterogeneous inputs, enabling precise and generalizable motor control (stepputtis2020language; karamcheti2023language; livision). Formally, we assume access to a dataset $\mathcal{D}=\{l,\tau_{i}\}_{i=1}^{N}$ of NN expert demonstrations, where each demonstration has a sequence of state-action pairs over horizon $H$, $\tau=\{(s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{H},a_{H})\}$, aligned with a task goal $l\in L$ (e.g., “slide the red cube”). The states $s_{t}=(o_{t},r_{t})\in S$ integrate current visual observation $o_{t}$ and robot proprioceptive data $r_{t}$, while actions $a_{t}\in A$ represent motor commands like end-effector positions or joint velocities. Our goal is to learn a policy $\pi_{\theta}:S\times L\rightarrow A$, which maps states $S$ and natural language instructions $L$ to actions AA. We optimize this policy $\pi_{\theta}$ by minimizing the negative log-likelihood of the observed expert actions:
@@
Metrics and Implementation Details. The success rate (%) measures task completion under defined conditions, such as transferring a red cube between grippers with 1cm clearance to avoid collisions. To ensure consistency, our model maintains the same settings as AWE (shiwaypoint; ke20243d) for Aloha MuJoCo and RLBench, and aligns with MPI (zeng2024learning) for the Franka Kitchen setup. The model employs the SGD optimizer, starting with a learning rate of 1e-5, momentum of 0.9, a chunking size of k=k=50, and a batch size of 88. The ODEsolver evaluates solutions at two discrete time points.
@@ Conclusion
We present CCoL, a novel framework that synergizes multi-modal continuous co-learning and stepwise semantic-physical alignment to mitigate compounding errors in BC. By integrating NeuralODEs for smooth latent trajectory learning and bidirectional cross-attention for linguistic grounding, CCoL ensures both temporal coherence and semantic precision during task execution. Extensive simulation experiments, alongside real-world robustness, demonstrate its superiority, especially in bimanual collaborative and long-horizon manipulation tasks. Future work will extend CCoL to LLM-based methods and explore its integration with foundation models for open-world manipulation.

### v4

@@ Introduction
  • We propose a multimodal continuous co-learning mechanism integrating vision, language, and proprioception to model temporal dependencies in latent space, ensuring smooth and consistent transitions across action states.
@@
LCM leverages BC methods to infer executable behaviors from heterogeneous inputs, enabling precise and generalizable motor control (stepputtis2020language; karamcheti2023language; livision). Formally, we assume access to a dataset $\mathcal{D}=\{l,\tau_{i}\}_{i=1}^{N}$ of $N$ expert demonstrations, where each demonstration has a sequence of state-action pairs over horizon $H$, $\tau=\{(s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{H},a_{H})\}$, aligned with a task goal $l\in L$ (e.g., “slide the red cube”). The states $s_{t}=(o_{t},r_{t})\in S$ integrate current visual observation $o_{t}$ and robot proprioceptive data $r_{t}$, while actions $a_{t}\in A$ represent motor commands like end-effector positions or joint velocities. Our goal is to learn a policy $\pi_{\theta}:S\times L\rightarrow A$, which maps states $S$ and natural language instructions $L$ to actions $A$. We optimize this policy $\pi_{\theta}$ by minimizing the negative log-likelihood of the observed expert actions:
@@
Metrics and Implementation Details. The success rate (%) measures task completion under defined conditions, such as transferring a red cube between grippers with 1cm clearance to avoid collisions. To ensure consistency, our model maintains the same settings as AWE (shiwaypoint; ke20243d) for Aloha MuJoCo and RLBench, and aligns with MPI (zeng2024learning) for the Franka Kitchen setup. The model employs the SGD optimizer, starting with a learning rate of 1e-5, momentum of 0.9, a chunking size of k=k=50, and a batch size of $8$. The ODEsolver evaluates solutions at two discrete time points.
@@ Conclusion
We present CCoL, a novel framework that synergizes multimodal continuous co-learning and stepwise semantic-physical alignment to mitigate compounding errors in BC. By integrating NeuralODEs for smooth latent trajectory learning and bidirectional cross-attention for linguistic grounding, CCoL ensures both temporal coherence and semantic precision during task execution. Extensive simulation experiments, alongside real-world robustness, demonstrate its superiority, especially in bimanual collaborative and long-horizon manipulation tasks. Future work will extend CCoL to LLM-based methods and explore its integration with foundation models for open-world manipulation.